OpenStudy (anonymous):

does (n^3)/(5n^(10/3)) = 1/5n^(1/3)?? Thanks in advance

OpenStudy (lalaly):

\[\frac{1}{5} \frac{n^3}{n^\frac{10}{3}} = \frac{1}{5}n^{-\frac{1}{3}} = \frac{1}{5n^{\frac{1}{3}}}\]

OpenStudy (lalaly):

\[=\frac{1}{5\sqrt[3]{n}}\]